OpenStudy (anonymous):

What number replaces the ? 12 = 12 12 + 1 + 4 = 12 + A. 1 B. 4 C. 5 D. 17

OpenStudy (anonymous):

well what is 12 + 1 + 4

OpenStudy (anonymous):

17

OpenStudy (anonymous):

and what is 17 - 12

OpenStudy (anonymous):

5

OpenStudy (anonymous):

there's your answer ^.^

OpenStudy (anonymous):

can u help me with more please

OpenStudy (anonymous):

but not the direct anmswers

OpenStudy (anonymous):

yes of course

OpenStudy (tkhunny):

Seriously? 12 + 1 + 4 = 12 + (1 + 4) = 12 + 5 = Why go to all that trouble?

OpenStudy (anonymous):

what trouble? I helped her just fine...

OpenStudy (anonymous):

there is not trouble

OpenStudy (anonymous):

Let p = 4. What is the value of q in this equation? 9 + p × 3 = q

OpenStudy (anonymous):

i got 39

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

yes I believe that is correct :)

OpenStudy (anonymous):

thx for all the help u guys are amazing and thank u autumn the most

OpenStudy (anonymous):

you are totally welcome ^.^

OpenStudy (anonymous):

aww

OpenStudy (tkhunny):

Whoops! 9 + p × 3 = q p = 4 9 + 4 × 3 = q Let's remember our Order of Operations 9 + (4 × 3) = q 9 + 12 = q How do you now feel about 39?